# Off the Planckian Locus: Using 2D Chromaticity to Improve In-Camera Color

### v2

@@ Abstract
Traditional in-camera colorimetric mapping relies on correlated color temperature (CCT)–based interpolation between pre-calibrated transforms optimized for Planckian illuminants such as CIE A and D65. However, modern lighting technologies such as LEDs can deviate substantially from the Planckian locus, exposing the limitations of relying on conventional one-dimensional CCT for illumination characterization. This paper demonstrates that transitioning from 1D CCT (on the Planckian locus) to a 2D chromaticity space (off the Planckian locus) improves colorimetric accuracy across various mapping approaches. In addition, we replace conventional CCT interpolation with a lightweight multi-layer perceptron (MLP) that leverages 2D chromaticity features for robust colorimetric mapping under non-Planckian illuminants. A lightbox-based calibration procedure incorporating representative LED sources is used to train our MLP. Validated across diverse LED lighting, our method reduces angular reproduction error by 22% on average in LED-lit scenes, maintains backward compatibility with traditional illuminants, accommodates multi-illuminant scenes, and supports real-time in-camera deployment with negligible additional computational cost. Project page: https://cst-mlp.github.io
@@ 1 Introduction
Figure 1: 1D Planckian interpolation vs. our 2D chromaticity-based prediction. Traditional 1D methods perform adequately for light sources whose CCTs lie close to the Planckian locus but are notably less accurate for light sources (namely, LEDs) that lie off the Planckian locus. Our approach leverages 2D chromaticity coordinates for more robust color reproduction across all illuminant types. Heat maps show angular reproduction error, from $0^{\circ}$ (dark blue) to $15^{\circ}$ (dark red), demonstrating comparable performance along the Planckian curve but substantial improvements in LED-dominated regions where existing methods break down.
@@
The current in-camera colorimetric mapping method was developed when early digital cameras primarily captured images illuminated by conventional light sources (e.g., incandescent, fluorescent, and natural/outdoor light). The perceived colors of these early light sources’ spectra are conveniently correlated with blackbody radiators—idealized physical bodies whose spectral distributions depend solely on temperature according to Planck’s law [wyszecki1982color]. In the CIE xy chromaticity space [xyz], these blackbody sources form the Planckian locus, as shown in Figure 1. Correlated color temperature (CCT) provides a simple 1D representation for characterizing such conventional light sources. The standard procedure for CST computation involves one-dimensional CCT-based interpolation of pre-calibrated matrices along the Planckian locus. Typically, two pre-calibrated illuminants are used (1) daylight 6500 K temperature and (2) incandescent light (CIE A).
However, the proliferation of LED technology has fundamentally disrupted this framework. Modern artificial illuminants exhibit chromaticity coordinates that deviate significantly from the Planckian locus, resulting in suboptimal color reproduction when using CCT-based interpolation (Figure 1). This mismatch highlights the growing disparity between current camera color-processing assumptions and the spectral characteristics of modern illumination.
This paper proposes shifting from a one-dimensional CCT space to a 2D chromaticity space for CST estimation. Our approach recognizes that modern illuminants require characterization beyond simple CCT, incorporating camera-specific chromaticity coordinates as input features. Moreover, we employ a lightweight multi-layered perceptron network to learn optimal interpolation. By training on an expanded dataset including representative LED illuminants alongside traditional sources, our CST-MLP achieves superior colorimetric mapping that adapts to contemporary lighting spectral diversity while maintaining backward compatibility.
@@ 2 Motivation and Related Work
Motivation. The colorimetric mapping stage of the digital camera pipeline (top/middle of Figure 2 consists of two main steps: white balance (WB) via a diagonal correction and color balance (CB) via a CST. In the camera pipeline, WB is typically modeled in the 2D raw chromaticity space [Hakki_eccv16]. Other software, such as Lightroom [lightroom], allows 2D WB control via temperature and tint [Ohno] sliders. While estimated WB values are represented in 2D chromaticity space, the color-balancing stage of the camera pipeline projects 2D chromaticities onto a 1D CCT and utilizes a CCT-based interpolation [AdobeDNG] to derive the CST (see Supplementary Section S5 for details on this procedure).
@@
When illumination chromaticity coordinates fall far from the Planckian locus, CCT calculations become ill-defined [Ohno], and subsequent CCT-based interpolation leads to errors exceeding just-noticeable color differences [Buzzelli]. Karaimer and Brown [Hakki_CVPR18] introduced an additional calibration point along the Planckian locus to improve interpolation, an approach later integrated into the Adobe DNG specification [AdobeDNG]. However, because it still relies on CCT-based interpolation, this method cannot represent light sources that fall off the Planckian locus. Other research has explored improved illumination estimation [gcc, Zhan, Dongyoung_2024, Juarez] and post-processing strategies [abc_chiu, deepwb, Finlayson, Hakki_CVPR18], yet these approaches do not address the fundamental limitation of CCT-based interpolation in the colorimetric mapping stage. To overcome this, we propose predicting the CST directly in the 2D chromaticity space using a CST-MLP, as illustrated in Figure 2 (bottom).
Figure 2: Top: Camera pipeline overview. Early stages convert raw sensor responses to colorimetric values, while later stages transform colors to a photo-finished display-referred space [Hakki_eccv16]. This paper aims to achieve maximum accuracy in the colorimetric stage. Middle: Colorimetric mapping stage. When an image is captured under arbitrary illumination, the estimated illumination is used to remove color cast via a diagonal white-balance $3\!\times\!3$ matrix (i.e., $W^{l_{a}}_{D}$), then mapped to CIE XYZ with an illumination-specific $3\!\times\!3$ CST matrix (i.e., $T_{l_{a}}$). Bottom: 2D chromaticity CST prediction vs. 1D CCT interpolation. CST interpolation approaches (e.g., $T_{l_{a}}=gT_{2500K}+(1-g)T_{6500K}$) assume CST parameters can be derived from a CCT-based interpolation, given calibrated illuminants along the Planckian Locus. CCT-based interpolation works poorly for non-Planckian illuminants. We propose a 2D CST-MLP that takes chromaticity coordinates as input and directly predicts the full $3\!\times\!3$ CST matrix.
@@
Figure 3: Our 2D chromaticity CST prediction framework. (Stage 1) Controlled lightbox dataset capture: A ColorChecker [xrite] target is captured under diverse illumination, including traditional Planckian sources and modern LED illuminants with varying chromaticity coordinates, ensuring comprehensive coverage of on-locus and off-locus illuminants. (Stage 2) CST-MLP training: Our lightweight MLP takes 2D chromaticity coordinates as input and predicts the color space transform directly, bypassing CCT-based interpolation. To train CST-MLP, we begin with raw color patches $P_{\text{raw}}$ and apply white balancing using the estimated illuminant $w$ (derived from the white patch), resulting in white-balanced patches $P_{\text{wb}}$. We then convert the illuminant $w$ into chromaticity coordinates $w_{xy}$ via the DNG optimization procedure (see Supplementary Section S5). The chromaticity $w_{xy}$ serves as input to our CST-MLP, which predicts the CST applied to $P_{\text{wb}}$. Finally, we train with a cosine loss between the corrected patches $P_{c}$ and the ground-truth patches $P_{\text{gt}}$ in XYZ space.
@@
Color balance.  Additional correction is required for non-neutral colors. Thus, diagonal correction provides white balance (WB), while full correction is color balance (CB). Diagonal correction suffices for particular illuminations (e.g., broadband) to achieve full CB, but many illuminations require full color balancing [cheng]. Camera ISPs apply an illumination-specific CST [Hakki_CVPR18] after WB to correct non-neutral colors. Adobe DNG implements this by mapping illuminants to CCT, then interpolating between calibrated $3\times 3$ CST matrices that map to CIE XYZ. Initially, DNG supported two CST matrices at high ($\sim$6500 K) and low ($\sim$2500 K) CCT. Karaimer and Brown [Hakki_CVPR18] showed that an additional calibration point ($\sim$4300 K) improves reproduction, resulting in DNG now supporting three CSTs. Other higher-dimensional mappings, such as polynomial [cheung] and root-polynomial [Finlayson], could be used in place of the $3\times 3$ transformation; however, these mappings are prone to overfitting [Finlayson] and are non-invertible.
@@
Direct mapping. Recent methods such as CCCNN [cccnn] and EXPINV [expinv] learn an illuminant-dependent colorimetric mapping by training MLPs that take raw pixel values and illuminant CCT as input to predict CIE XYZ values directly. However, these networks require expensive per-pixel evaluation at test time and rely on access to camera spectral sensitivities during training. Their pixel-wise formulation introduces spatial inconsistencies and color casts, as we demonstrate in Section 4.2. Finally, their dependence on 1D CCT illumination coordinates limits these networks’ ability to model non-Planckian illuminants. In contrast, our proposed CST-MLP predicts a global CST in a single forward pass, enforces spatial consistency, requires no spectral sensitivity data, and generalizes better across diverse light sources by operating in 2D chromaticity space.
@@ 3 Method
We aim to learn a mapping from illuminant chromaticity to a CST using a two-stage method, visualized in Figure 3, involving calibration data capture within a lightbox (Section 3.1) and training a lightweight CST-MLP (Section 3.2).
@@ 3.1 Data Capture
Traditional CST calibration methods calibrate CSTs by capturing color charts under standard illuminants such as CIE A or D65. However, we require our method to handle non-Planckian illuminants that a CCT cannot adequately describe. The first step in our process is to capture calibration data across diverse illuminants so our model learns the correct CST under diverse lighting conditions.
@@
We capture each color chart under 790 different illuminant SPDs sourced from two methods. First, we utilize 390 illuminants from Punnappurath et al. [Hoang] that estimate various light sources, including sunlight, fluorescent, incandescent, and LED. We use the original splits of 195 training, 78 validation, and 117 test illuminants. Then, we introduce a set of 400 illuminants generated by linearly combining the Telelumen LED SPDs with weights randomly sampled from a Dirichlet distribution (see Supplementary Section S2 for details). Using this distribution results in an even spread of illuminants across the chromaticity space and allows us to sample more illuminants farther from the Planckian Locus. We randomly split the dataset into 200 train, 80 validation, and 120 test illuminants. We combine both datasets for our calibration, resulting in 395 train, 158 validation, and 237 test illuminants.
@@ 3.2 Model
We propose a model, CST-MLP, that maps a white-point chromaticity $w_{xy}$ to a CST $T$. Following Karaimer and Brown [Hakki_CVPR18], we use the setting of mapping white-balanced raw images into the XYZ canonical space, and assume the raw image $I_{raw}$ and ground-truth white point chromaticity $w=[r/g,b/g]$ are known. Following previous work [Hakki_CVPR18], we use the ground-truth white point from the color chart white patch and compute the white-balanced raw image $I_{wb}$, by applying a diagonal correction [barron2015convolutional]. For our model, we convert the raw-space chromaticity $w$ to its xy-space counterpart $w_{xy}$ using the DNG [AdobeDNG] optimization procedure described in Supplementary Section S5.
CST-MLP is a small network with one hidden layer of 32 units that predicts the eight free entries of a $3\times 3$ CST matrix. Since the CST is defined only up to scale due to the angular-error loss, we fix the central element of the matrix to $1$ to remove this ambiguity. The eight predicted values are then reshaped into the CST matrix. CST-MLP is constrained to predict linear color transformations to maintain compatibility with the DNG pipeline. Polynomial and root-polynomial variants are discussed in Supplementary Section S1.3.
@@ 4 Evaluation
First, we evaluate 1D CCT vs. 2D chromaticity-based approaches using the light box dataset described in Section 3.1. Next, we test our light-box calibration across a set of in-the-wild scenes. Finally, we compare our method on single-illuminant scenes from the LSMI [lsmi] dataset and further extend the evaluation to multi-illuminant settings.
@@ 4.1 Evaluation Details
For comparison, we implement a wide variety of state-of-the-art baselines. First, to contextualize the upper bound of CST prediction, we compute the Oracle, the best-fit $3\times 3$ matrix that minimizes the angular error given the full color chart. Then, we compare with the standard 2-CST interpolation described by the DNG [AdobeDNG], which uses matrices recalibrated by capturing the color chart under Standard Illuminant A (2850 K) and D65 (6500 K) lighting within the lightbox. We also compare to the 3-CST [Hakki_CVPR18] interpolation with an additional calibrated CST at 4250 K. Next, we implement a nearest neighbors (NN) [nearestneighbors] baseline that maps each training white point with its best-fit $3\times 3$ CST (minimizing angular error on the color chart). At test time, given a white-point coordinate, we retrieve the nearest illuminant by L2 distance and apply its corresponding matrix. Finally, we compare direct MLP methods, CCCNN [cccnn] and EXPINV [expinv], that perform per-pixel correction given an input color and white point.
1D CCT vs 2D Chromaticity.
@@
To demonstrate the importance of using 2D chromaticity coordinates, we also compare our CST-MLP with a 1D CCT-based variant. Additionally, we adapt the 1D CCT-based baseline methods (NN, EXPINV, and CCCNN) into 2D chromaticity-based versions to show the benefits of our approach.
Training details.
@@
We use the dataset splits in Section 3.1 to train and evaluate all methods outside of CCCNN [cccnn] and EXPINV [expinv]. As described in the original papers, CCCNN [cccnn] and EXPINV [expinv] require a larger simulated dataset, as these networks aim to learn a per-pixel mapping. For simulation, we measure each camera’s spectral sensitivities using CamspecV2 [camspec], illumination SPDs from our lightbox capture setup (see Section 3.1), and 1269 reflectance SPDs from the Munsell Book of Color [munsell]. We then synthesize data using the image formation model described in CCCNN [cccnn] and EXPINV [expinv].
@@
All MLP methods are trained for 100,000 iterations using the Adam [adam] optimizer with a learning rate of 0.001. Our CST-MLP is trained using the cosine loss described in Section 3.2, while CCCNN [cccnn] and EXPINV [expinv] are trained with the recommended $\Delta E$ loss.
Noise augmentation.
When queried with out-of-domain chromaticities not seen in our training set, our 2D CST-MLP can sometimes fail to generalize. We fix this issue by adding Gaussian noise to input chromaticity coordinates at training time [ling2025stochastic], allowing models to observe chromaticities that may be out of distribution. For CST-MLP, we add noise with standard deviation $\sigma_{0}=0.05$ to the normalized input coordinates for all training iterations. We show the positive effects of this augmentation in our ablation study (Table 6). Finally, we exclude this augmentation for direct-prediction MLP baselines (EXPINV [expinv], CCCNN [cccnn]) as it can degrade their performance (see Supplementary Section S1.4).
Metrics.
For evaluation, we utilize angular error following previous work [Hakki_CVPR18]. We also compare $\Delta E_{2000}$ after applying a uniform scaling to all colors to match the Y channel of the third gray patch (bottom row of color chart). Finally, we report model parameter sizes (in KB) and inference costs (in MACs) when applied to a $1080\times 720$ image.
4.2 Lightbox Dataset Results
We compare all methods quantitatively on our dataset in Table 1 and qualitatively in Figure 4 (additional visualizations in Supplementary Section S2). We find that all methods benefit from using a 2D chromaticity coordinate instead of a 1D CCT, which confirms our motivation for this work. Additionally, our 2D CST-MLP surpasses traditional interpolation and nearest-neighbor methods, providing a strong solution that integrates seamlessly into the existing DNG [AdobeDNG] pipeline, which relies on linear transformations.
@@ 4.2 Lightbox Dataset Results
Mean 25% 50% 75% 90% Mean 25% 50% 75% 90% Mean 25% 50% 75% 90% Mean 25% 50% 75% 90%
Oracle 1D 0.00 0.01 1.71 0.54 1.15 2.14 3.68 2.06 0.64 1.31 2.63 4.65 2.79 0.95 2.00 3.49 5.94 2.58 0.84 1.83 3.25 5.47
2-CST 1D 0.07 0.01 3.31 0.80 1.98 4.33 7.91 3.86 0.96 2.31 5.38 9.59 4.56 1.48 3.25 6.03 9.94 4.51 1.53 3.19 5.93 9.95
3-CST 1D 0.11 0.01 3.31 0.80 2.01 4.31 7.91 3.86 0.94 2.32 5.39 9.68 4.53 1.48 3.22 5.97 9.95 4.51 1.51 3.18 5.92 9.94
NN 1D 15.43 0.01 3.93 0.94 2.33 5.21 9.57 4.68 1.06 2.75 6.43 11.57 5.31 1.54 3.64 7.32 11.84 5.46 1.86 3.54 7.21 12.36
CCCNN 1D 68.51 13.39 3.13 0.99 2.02 4.07 7.21 3.58 1.13 2.40 4.71 8.14 4.41 1.53 3.10 5.81 9.64 4.63 1.53 3.34 6.25 10.05
EXPINV 1D 68.53 13.39 3.14 0.90 1.92 4.06 7.57 3.62 0.96 2.29 4.84 8.59 4.29 1.35 2.82 5.74 9.65 4.69 1.45 3.43 6.35 10.25
CST-MLP 1D 1.41 0.01 3.00 0.70 1.63 3.76 7.54 3.42 0.67 1.85 4.50 8.88 4.03 1.05 2.48 5.31 9.74 4.00 1.10 2.52 5.19 9.54
NN 2D 16.97 0.01 3.14 0.82 1.79 3.87 7.47 3.54 0.90 2.04 4.61 8.50 4.25 1.39 2.87 5.22 9.57 4.15 1.33 2.72 5.34 9.48
CCCNN 2D 69.01 13.48 2.89 0.97 1.99 3.83 6.43 3.00 1.13 2.20 3.91 6.19 4.00 1.50 2.93 5.20 8.76 4.17 1.78 3.32 5.51 8.52
EXPINV 2D 69.03 13.48 2.66 0.87 1.78 3.45 5.97 3.11 0.94 2.26 4.16 6.89 3.84 1.32 2.71 5.02 8.41 4.14 1.49 3.05 5.44 8.86
CST-MLP 2D 1.54 0.01 2.58 0.66 1.52 3.36 6.22 2.85 0.68 1.69 3.70 6.78 3.57 1.03 2.30 4.63 8.34 3.61 1.05 2.34 4.62 8.34

### v3

@@ Abstract
Traditional in-camera colorimetric mapping relies on correlated color temperature (CCT)–based interpolation between pre-calibrated transforms optimized for Planckian illuminants such as CIE A and D65. However, modern lighting technologies such as LEDs can deviate substantially from the Planckian locus, exposing the limitations of relying on conventional one-dimensional CCT for illumination characterization. This paper demonstrates that transitioning from 1D CCT (on the Planckian locus) to a 2D chromaticity space (off the Planckian locus) improves colorimetric accuracy across various mapping approaches. In addition, we replace conventional CCT interpolation with a lightweight multi-layer perceptron (MLP) that leverages 2D chromaticity features for robust colorimetric mapping under non-Planckian illuminants. A lightbox-based calibration procedure incorporating representative LED sources is used to train our MLP. Validated across diverse LED lighting, our method reduces angular reproduction error by 22% on average in LED-lit scenes, maintains backward compatibility with traditional illuminants, accommodates multi-illuminant scenes, and supports real-time in-camera deployment with negligible additional computational cost.
@@ 1 Introduction
Figure 1: 1D Planckian interpolation vs. our 2D chromaticity-based prediction. Traditional 1D methods perform adequately for light sources whose CCTs lie close to the Planckian locus but are less accurate for light sources (namely, LEDs) that lie off the Planckian locus. Our approach leverages 2D chromaticity coordinates for more robust color reproduction across all illuminant types. Heat maps show angular reproduction error, from $0^{\circ}$ (dark blue) to $15^{\circ}$ (dark red), demonstrating comparable performance along the Planckian curve but substantial improvements in LED-dominated regions where existing methods break down.
@@
The current in-camera colorimetric mapping method was developed when early digital cameras primarily captured images illuminated by conventional light sources (e.g., incandescent, fluorescent, and natural/outdoor light). The perceived colors of these early light sources’ spectra are conveniently correlated with blackbody radiators—idealized physical bodies whose spectral distributions depend solely on temperature according to Planck’s law [wyszecki1982color]. In the CIE xy chromaticity space [xyz], these blackbody sources form the Planckian locus, as shown in Fig.˜1. Correlated color temperature (CCT) provides a simple 1D representation for characterizing such conventional light sources. The standard procedure for CST computation involves one-dimensional CCT-based interpolation of pre-calibrated matrices along the Planckian locus. Typically, two pre-calibrated illuminants are used (1) daylight (6500 K) and (2) incandescent light (CIE A).
However, the proliferation of LED technology has fundamentally disrupted this framework. Modern artificial illuminants exhibit chromaticity coordinates that deviate significantly from the Planckian locus, resulting in suboptimal color reproduction when using CCT-based interpolation (Fig.˜1). This mismatch highlights the growing disparity between current camera color-processing assumptions and the spectral characteristics of modern illumination.
This paper proposes shifting from a one-dimensional CCT space to a 2D chromaticity space for CST estimation. Our approach recognizes that modern illuminants require characterization beyond simple CCT, incorporating camera-specific chromaticity coordinates as input features. Moreover, we employ a low-parameter multi-layer perceptron (MLP) to learn optimal interpolation. By training on an expanded dataset including representative LED illuminants alongside traditional sources, our CST-MLP achieves superior colorimetric mapping that adapts to contemporary lighting spectral diversity while maintaining backward compatibility.
@@ 2 Motivation and Related Work
Motivation. The colorimetric mapping stage of the digital camera pipeline (top/middle of Fig.˜2 consists of two main steps: white balance (WB) via a diagonal correction and color balance (CB) via a CST. In the camera pipeline, WB is typically modeled in the 2D raw chromaticity space [Hakki_eccv16]. Other software, such as Lightroom [lightroom], allows 2D WB control via temperature and tint [Ohno] sliders. While estimated WB values are represented in 2D chromaticity space, the color-balancing stage of the camera pipeline projects 2D chromaticities onto a 1D CCT and utilizes a CCT-based interpolation [AdobeDNG] to derive the CST (see Supplementary Sec.˜S5 for details on this procedure).
@@
When illumination chromaticity coordinates fall far from the Planckian locus, CCT calculations become ill-defined [Ohno], and subsequent CCT-based interpolation leads to errors exceeding just-noticeable color differences [Buzzelli]. Karaimer and Brown [Hakki_CVPR18] introduced an additional calibration point along the Planckian locus to improve interpolation, an approach later integrated into the Adobe DNG specification [AdobeDNG]. However, because it still relies on CCT-based interpolation, this method cannot represent light sources that fall off the Planckian locus. Other research has explored improved illumination estimation [gcc, Zhan, Dongyoung_2024, Juarez] and post-processing strategies [abc_chiu, deepwb, Finlayson, Hakki_CVPR18], yet these approaches do not address the fundamental limitation of CCT-based interpolation in the colorimetric mapping stage. To overcome this, we propose predicting the CST directly in the 2D chromaticity space using a CST-MLP, as illustrated in Fig.˜2 (bottom).
Figure 2: Top Left: Camera pipeline overview. Early stages convert raw sensor responses to colorimetric values, while later stages transform colors to a photo-finished display-referred space [Hakki_eccv16]. This paper aims to achieve maximum accuracy in the colorimetric stage. Bottom Left: Colorimetric mapping stage. When an image is captured under arbitrary illumination, the estimated illumination is used to remove color cast via a diagonal white-balance $3\!\times\!3$ matrix (i.e., $W^{l_{a}}_{D}$), then mapped to CIE XYZ with an illumination-specific $3\!\times\!3$ CST matrix (i.e., $T_{l_{a}}$). Right: 2D chromaticity CST prediction vs. 1D CCT interpolation. CST interpolation approaches (e.g., $T_{l_{a}}=gT_{2500K}+(1-g)T_{6500K}$) assume CST parameters can be derived from a CCT-based interpolation, given calibrated illuminants along the Planckian Locus. CCT-based interpolation works poorly for non-Planckian illuminants. We propose a 2D CST-MLP that takes chromaticity coordinates as input and directly predicts the full $3\!\times\!3$ CST matrix.
@@
Figure 3: Our 2D chromaticity CST prediction framework. (Stage 1) Controlled lightbox dataset capture: A ColorChecker [xrite] target is captured under diverse illumination, including traditional Planckian sources and modern LED illuminants with varying chromaticity coordinates, ensuring comprehensive coverage of on-locus and off-locus illuminants. (Stage 2) CST-MLP training: Our lightweight MLP takes 2D chromaticity coordinates as input and predicts the color space transform directly, bypassing CCT-based interpolation. To train CST-MLP, we begin with raw color patches $P_{\text{raw}}$ and apply white balancing using the estimated illuminant $w$ (derived from the white patch), resulting in white-balanced patches $P_{\text{wb}}$. We then convert the illuminant $w$ into chromaticity coordinates $w_{xy}$ via the DNG optimization procedure (see Supplementary Sec.˜S5). The chromaticity $w_{xy}$ serves as input to our CST-MLP, which predicts the CST applied to $P_{\text{wb}}$. Finally, we train with a cosine loss between the corrected patches $P_{c}$ and the ground-truth patches $P_{\text{gt}}$ in XYZ space.
@@
Color balance.  Additional correction is required for non-neutral colors. Thus, diagonal correction provides white balance (WB), while full correction is color balance (CB). Diagonal correction suffices for particular illuminations (e.g., broadband) to achieve full CB, but many illuminations require full color balancing [cheng]. Camera ISPs apply an illumination-specific CST [Hakki_CVPR18] after WB to correct non-neutral colors. Adobe DNG implements this by mapping illuminants to CCT, then interpolating between calibrated $3\times 3$ CST matrices that map to CIE XYZ. Initially, DNG supported two CST matrices at high ($\sim$6500 K) and low ($\sim$2500 K) CCT. Karaimer and Brown [Hakki_CVPR18] showed that an additional calibration point ($\sim$4300 K) improves reproduction, resulting in DNG now supporting three CSTs. Other higher-dimensional mappings, such as polynomial [cheung] and root-polynomial [Finlayson], could be used in place of the $3\times 3$ transformation when invertibility is not required.
@@
Direct mapping. Recent methods such as CCCNN [cccnn] and EXPINV [expinv] learn an illuminant-dependent colorimetric mapping by training MLPs that take raw pixel values and illuminant CCT as input to predict CIE XYZ values directly. However, these networks require expensive per-pixel evaluation at test time and rely on access to camera spectral sensitivities during training. Their pixel-wise formulation introduces spatial inconsistencies and color casts, as we demonstrate in Sec.˜4.2. Finally, their dependence on 1D CCT illumination coordinates limits these networks’ ability to model non-Planckian illuminants. In contrast, our proposed CST-MLP predicts a global CST in a single forward pass, enforces spatial consistency, requires no spectral sensitivity data, and generalizes better across diverse light sources by operating in 2D chromaticity space.
@@ 3 Method
We aim to learn a mapping from illuminant chromaticity to a CST using a two-stage method, visualized in Fig.˜3, involving calibration data capture within a lightbox (Sec.˜3.1) and training a lightweight CST-MLP (Sec.˜3.2).
@@ 3.1 Data Capture
Traditional CST calibration methods calibrate CSTs by capturing color charts under standard illuminants such as CIE A or D65. However, we require our method to handle non-Planckian illuminants that a CCT cannot adequately describe. The first step in our process is capturing calibration data across diverse illuminants so our model learns the correct CST under diverse lighting conditions.
@@
We capture each color chart under 790 different illuminant SPDs sourced from two methods. First, we utilize 390 illuminants from Punnappurath et al. [Hoang] that estimate various light sources, including sunlight, fluorescent, incandescent, and LED. We use the original splits of 195 training, 78 validation, and 117 test illuminants. Then, we introduce a set of 400 illuminants generated by linearly combining the Telelumen LED SPDs with weights randomly sampled from a Dirichlet distribution (see Supplementary Sec.˜S2 for details). Using this distribution results in an even spread of illuminants across the chromaticity space and allows us to sample more illuminants farther from the Planckian Locus. We randomly split the dataset into 200 train, 80 validation, and 120 test illuminants. We combine both datasets for our calibration, resulting in 395 train, 158 validation, and 237 test illuminants.
@@ 3.2 Model
We propose a model, CST-MLP, that maps a white-point chromaticity $w_{xy}$ to a CST $T$. Following Karaimer and Brown [Hakki_CVPR18], we use the setting of mapping white-balanced raw images into the XYZ canonical space, and assume the raw image $I_{raw}$ and ground-truth white point chromaticity $w=[r/g,b/g]$ are known. Following previous work [Hakki_CVPR18], we use the ground-truth white point from the color chart white patch and compute the white-balanced raw image $I_{wb}$, by applying a diagonal correction [barron2015convolutional]. For our model, we convert the raw-space chromaticity $w$ to its xy-space counterpart $w_{xy}$ using the DNG [AdobeDNG] optimization procedure described in Supplementary Sec.˜S5. We note that this optimization is a fast, commonly implemented algorithm [Hakki_CVPR18] on modern phones and executes in an average of $250\ \mu s$ on an iPhone 16.
CST-MLP is a small network with one hidden layer of 32 units that predicts the eight free entries of a $3\times 3$ CST matrix. Since the CST is defined only up to scale due to the angular-error loss, we fix the central element of the matrix to $1$ to remove this ambiguity. The eight predicted values are then reshaped into the CST matrix. If invertibility is not a requirement, we also discuss predicting polynomial and root-polynomial mappings in Sec.˜4.
@@ 4 Evaluation
First, we evaluate 1D CCT vs. 2D chromaticity-based approaches using the light box dataset described in Sec.˜3.1. Next, we test our light-box calibration across a set of in-the-wild scenes. Finally, we compare our method on single-illuminant scenes from the LSMI [lsmi] dataset and further extend the evaluation to multi-illuminant settings.
@@ 4.1 Evaluation Details
Baselines. For comparison, we implement a variety of state-of-the-art baselines. First, to contextualize the upper bound of CST prediction, we compute the Oracle, the best-fit $3\times 3$ matrix that minimizes the angular error given the full color chart. Then, we compare with the standard 2-CST interpolation described by the DNG [AdobeDNG], which uses matrices recalibrated by capturing the color chart under Standard Illuminant A (2850 K) and D65 (6500 K) lighting within the lightbox. We also compare to the 3-CST [Hakki_CVPR18] interpolation with an additional calibrated CST at 4250 K. Next, we implement a nearest neighbors (NN) [nearestneighbors] baseline that maps each training white point with its best-fit $3\times 3$ CST (minimizing angular error on the color chart). At test time, given a white-point coordinate, we retrieve the nearest illuminant by L2 distance and apply its corresponding matrix. Finally, we compare direct MLP methods, CCCNN [cccnn] and EXPINV [expinv], that perform per-pixel correction given an input color and white-point.
@@
1D CCT vs 2D Chromaticity. To demonstrate the importance of using 2D chromaticity coordinates, we also compare our CST-MLP with a 1D CCT-based variant. Additionally, we adapt the 1D CCT-based baseline methods (NN, EXPINV, and CCCNN) into 2D chromaticity-based versions to show the benefits of our approach.
Polynomial mappings. To evaluate the effect of higher-dimensional transforms, we extend our CST-MLP to output polynomial [cheung] and root-polynomial [Finlayson] mappings. We compare standard linear $3\times 3$ transforms against polynomial and root-polynomial mappings of varying sizes and find that root-polynomial $3\times 6$ mappings yield the lowest angular error (see Tab.˜2). We denote the $3\times 6$ root-polynomial version as CST-MLP (RP) in all subsequent tables.
@@
Training details. We use the dataset splits in Sec.˜3.1 to train and evaluate all methods outside of CCCNN [cccnn] and EXPINV [expinv]. As described in the original papers, CCCNN [cccnn] and EXPINV [expinv] require a larger simulated dataset, as these networks aim to learn a per-pixel mapping. For simulation, we measure each camera’s spectral sensitivities using CamspecV2 [camspec], illumination SPDs from our lightbox capture setup (see  Sec.˜3.1), and 1269 reflectance SPDs from the Munsell Book of Color [munsell]. We then synthesize data using the image formation model described in CCCNN [cccnn] and EXPINV [expinv].
Figure 4: We visualize the chromaticities of our train (lightbox) and in-the-wild illuminants, which include a few out-of-distribution illuminants (circled purple).
@@
All MLP methods are trained for 100,000 iterations using the Adam [adam] optimizer with a learning rate of 0.001. Our CST-MLP is trained using the cosine loss described in Sec.˜3.2, while CCCNN [cccnn] and EXPINV [expinv] are trained with the recommended $\Delta E$ loss.
Table 1: Quantitative evaluation on lightbox dataset. We compare angular error and $\Delta E_{2000}$ of traditional 2-CST/3-CST interpolation [AdobeDNG, Hakki_CVPR18], NN [nearestneighbors], CCCNN [cccnn], EXPINV [expinv], and our CST-MLP. For NN, CCCNN, EXPINV, and CST-MLP, we evaluate both 1D CCT implementations and our proposed 2D chromaticity adaptations (highlighted in blue). Methods Marked (RP) predict a root-polynomial [Finlayson] mapping.
@@
Mean 25% 50% 90% Mean 25% 50% 90% Mean 25% 50% 90% Mean 25% 50% 90%
Oracle 1D 0.00 0.01 1.71 0.54 1.15 3.68 2.06 0.64 1.31 4.65 2.79 0.95 2.00 5.94 2.58 0.84 1.83 5.47
2-CST [AdobeDNG] 1D 0.07 0.01 3.31 0.80 1.98 7.91 3.86 0.96 2.31 9.59 4.56 1.48 3.25 9.94 4.51 1.53 3.19 9.95
3-CST [Hakki_CVPR18] 1D 0.11 0.01 3.31 0.80 2.01 7.91 3.86 0.94 2.32 9.68 4.53 1.48 3.22 9.95 4.51 1.51 3.18 9.94
NN [nearestneighbors] 1D 15.43 0.01 3.93 0.94 2.33 9.59 4.68 1.05 2.74 11.62 5.24 1.54 3.56 11.73 5.48 1.82 3.59 12.27
CCCNN [cccnn] 1D 68.51 13.39 3.28 1.09 2.13 7.38 3.58 1.19 2.42 8.02 4.39 1.47 2.96 9.67 4.71 1.76 3.35 10.00
EXPINV [expinv] 1D 68.53 13.39 3.13 0.91 1.86 7.42 3.61 1.08 2.33 8.46 4.31 1.26 2.85 9.88 4.68 1.43 3.31 10.40
CST-MLP 1D 1.41 0.01 3.09 0.91 1.85 7.35 3.52 0.98 2.06 8.60 4.11 1.37 2.62 9.55 4.05 1.29 2.60 9.42
CST-MLP(RP) 1D 2.57 0.01 2.82 0.85 1.73 6.65 3.26 0.96 2.05 7.68 3.77 1.34 2.45 8.57 3.70 1.21 2.43 8.52
NN [nearestneighbors] 2D 16.97 0.01 3.24 0.82 1.79 7.73 3.56 0.90 2.03 8.51 4.30 1.37 2.84 9.81 4.15 1.33 2.72 9.48
CCCNN [cccnn] 2D 69.01 13.48 2.89 0.97 1.99 6.43 3.00 1.13 2.20 6.19 4.00 1.50 2.93 8.76 4.17 1.78 3.32 8.52
EXPINV [expinv] 2D 69.03 13.48 2.66 0.87 1.78 5.97 3.11 0.94 2.26 6.89 3.84 1.32 2.71 8.41 4.14 1.49 3.05 8.86
CST-MLP 2D 1.54 0.01 2.66 0.82 1.70 6.14 2.89 0.87 1.85 6.67 3.60 1.22 2.38 8.18 3.60 1.20 2.43 8.23
CST-MLP(RP) 2D 2.70 0.01 2.36 0.71 1.44 5.53 2.60 0.71 1.58 6.17 3.27 1.13 2.14 7.47 3.28 1.17 2.20 7.25